OpenStudy (anonymous):

please help Given the following functions f(x) and g(x), solve (f + g)(3) and select the correct answer below: f(x) = 6x + 3 g(x) = x − 7

OpenStudy (anonymous):

@DanJS

OpenStudy (anonymous):

@DullJackel09

OpenStudy (anonymous):

@Donblue

OpenStudy (anonymous):

@queen-of-tokyo

jimthompson5910 (jim_thompson5910):

are you able to evaluate f(x) when x = 3 ?

OpenStudy (anonymous):

yes its 21 right?

OpenStudy (anonymous):

are u there?

jimthompson5910 (jim_thompson5910):

correct, f(3) = 21

jimthompson5910 (jim_thompson5910):

how about the value of g(3) ?

OpenStudy (anonymous):

-4?

jimthompson5910 (jim_thompson5910):

very good

jimthompson5910 (jim_thompson5910):

now just add the values of f(3) and g(3)

jimthompson5910 (jim_thompson5910):

this works because (f+g)(x) = f(x) + g(x) (f+g)(3) = f(3) + g(3)

OpenStudy (anonymous):

17 times 3?

jimthompson5910 (jim_thompson5910):

add

OpenStudy (anonymous):

4 17 25 31

jimthompson5910 (jim_thompson5910):

(f+g)(3) = f(3) + g(3) (f+g)(3) = 21 + (-4)

OpenStudy (anonymous):

those are the choices

jimthompson5910 (jim_thompson5910):

what do you mean by 17 times 3 ?

OpenStudy (anonymous):

because 21 +-4 is 17

jimthompson5910 (jim_thompson5910):

which is your answer

jimthompson5910 (jim_thompson5910):

I'm not sure why you are multiplying by 3

OpenStudy (anonymous):

because( f+g)(3

OpenStudy (anonymous):

u have to multiply by 3 after u add em right?

jimthompson5910 (jim_thompson5910):

the notation `(f+g)(3)` does NOT mean `(f+g) times 3`

jimthompson5910 (jim_thompson5910):

f(3) does not mean f times 3 it means "replace every x with 3, then evaluate"

OpenStudy (anonymous):

ok could u tell me what it means

OpenStudy (anonymous):

o ok

jimthompson5910 (jim_thompson5910):

(f+g)(3) is just another function that is evaluated at x = 3

OpenStudy (anonymous):

so its 17?

jimthompson5910 (jim_thompson5910):

yes, that is correct

OpenStudy (anonymous):

can u help with 1 more?

jimthompson5910 (jim_thompson5910):

sure

OpenStudy (anonymous):

ill create new thread then tagu

jimthompson5910 (jim_thompson5910):

ok